### v1

# SDEval: Safety Dynamic Evaluation for Multimodal Large Language Models

Abstract
In the rapidly evolving landscape of Multimodal Large Language Models (MLLMs), the safety concerns of their outputs have earned significant attention. Although numerous datasets have been proposed, they may become outdated with MLLM advancements and are susceptible to data contamination issues. To address these problems, we propose SDEval, the first safety dynamic evaluation framework to controllably adjust the distribution and complexity of safety benchmarks. Specifically, SDEval mainly adopts three dynamic strategies: text, image, and text-image dynamics to generate new samples from original benchmarks. We first explore the individual effects of text and image dynamics on model safety. Then, we find that injecting text dynamics into images can further impact safety, and conversely, injecting image dynamics into text also leads to safety risks. SDEval is general enough to be applied to various existing safety and even capability benchmarks. Experiments across safety benchmarks, MLLMGuard and VLSBench, and capability benchmarks, MMBench and MMVet, show that SDEval significantly influences safety evaluation, mitigates data contamination, and exposes safety limitations of MLLMs. Code is available at https://github.com/hq-King/SDEval
1 Introduction
Large language models (LLMs) (Brown et al. 2020; Achiam et al. 2023; Reid et al. 2024; DeepSeek-AI et al. 2024) have achieved significant advancements. Recent developments have extended this success into the multi-modal realm, allowing LLMs to execute various high-level vision tasks, including visual content understanding and generation (Xie et al. 2024; Li et al. 2024b; Gao et al. 2025; OpenAI 2024, 2025; Zhao et al. 2025). Despite the success in MLLM capabilities, there is a huge risk that MLLMs may generate outputs that diverge from their creators’ intended goals, potentially resulting in untruthful or harmful content (Hendrycks, Mazeika, and Woodside 2023; Yao et al. 2024). This highlights the crucial need for ensuring MLLM safety before deployment. Comprehensive assessment of their potential risks and corresponding mitigation strategies are needed.
Recently, several studies have initiated preliminary explorations into evaluating the safety of MLLMs. MLLMGuard (Gu et al. 2024) provides the safety analysis in both English and Chinese, using data from social media. Hu et al. (2024) identified information leakage issues in the existing datasets and proposed VLSBench, improving evaluation accuracy by better aligning image and text modalities. Besides, existing efforts also establish relatively comprehensive safety evaluation systems (Ying et al. 2024; Zhang et al. 2024; Cai et al. 2024; Liu et al. 2023b). However, after reviewing existing benchmarks, we identify the following main challenges in achieving reliable safety evaluation: 1) Data leakage. Most safety benchmarks build their dataset by integrating open-source datasets(Zhang et al. 2024; Gu et al. 2024; Zhang et al. 2025a; Xia et al. 2025; Liu et al. 2025), which are likely to be included in the MLLM training sets. Affected by this, the results of MLLMs on these benchmarks may lead to concerns, causing a misunderstanding in the entire community. 2) Static dataset with fixed complexity. Existing MLLM safety benchmarks are manually constructed and lack updating. Their fixed complexity can’t match the fast progress of MLLM. To gauge MLLM performance limits precisely, there’s an urgent need for a dynamic, automated evaluation framework with adjustable complexity. 3) Attack methods continue to evolve. As new attack methods emerge, MLLM safety benchmarks should be updated accordingly to further test model safety performance. Although previous studies have proposed simple dynamic evaluation methods (Yang et al. 2024b; Zhu et al. 2023), they are typically applied to model capability evaluation, neglecting safety and the balance of capability-safety, which hinders effective dynamic safety evaluation.
(a) Safety Rate
(b ) ASD and PAR
To tackle these challenges, we propose SDEval, a novel, general, and flexible framework for safety dynamic evaluation of MLLMs. To dynamically create new evaluation suites with flexible complexity, we divide the dynamic strategies into three parts: 1) Text Dynamics, which aims to figure out whether MLLMs can grasp the critical safety information in the prompt, which is presented in different types of expressions. We generate the new texts using methods such as character perturbation, linguistic mix, chain-of-thought injection, and so on. 2) Image Dynamics, which aims to explore whether MLLMs can consistently focus on safety-related subjects in images without being disturbed by other factors. We utilize tricks like diffusion-based generation and editing to modify original images. 3) Text-Image Dynamics, aiming to evaluate whether MLLMs can provide a deeper understanding of the safety of image-text pairs, and whether MLLMs can cope with common jailbreaking inputs. We focus on the combined impact of images and text on safety, as well as the influence of their interaction on safety. By integrating text and image dynamics into a comprehensive framework, SDEval can significantly improve data complexity and difficulty, as shown in Figure 3. SDEval is general and flexible, which can co-exist and co-evolve with existing benchmarks. Additionally, SDEval can also be utilized for capability dynamic evaluation. From a capability-safety balance perspective, SDEval reveals that most models exhibit greater instability in safety compared to capability, indicating an urgent requirement for further improvements in model safety.
We leverage SDEval for representative safety evaluation benchmarks such as MLLMGuard (Gu et al. 2024) and VLSBench (Hu et al. 2024), and capability evaluation benchmarks, MMVet (Yu et al. 2023) and MMBench (Liu et al. 2023d). Experiments on various MLLMs, e.g., GPT-4o (OpenAI 2024), Claude-4-Sonnet (Anthropic. 2025), and DeepSeek-VL family (Lu et al. 2024), demonstrate that SDEval impacts the safety of different MLLMs to varying degrees, with InternVL-3-78B experiencing a safety reduction of nearly 10%. These results indicate that our dynamic strategy significantly alleviates the data leakage problem, changes data distribution, and increases dataset complexity.
In summary, our contributions are the follows:
We proposed SDEval, the first safety dynamic evaluation framework for MLLMs. SDEval is general enough to be applied to various benchmarks and exhibits resistance to saturation for capability evaluation benchmarks.
We design a diverse set of text, image, and text-image interaction dynamic strategies, and conduct a detailed analysis of their dynamic effects.
We perform extensive experiments and ablation studies to validate the proposed strategy. Experiments demonstrate that our dynamic strategies effectively increase dataset complexity and reduce safety evaluation scores.
2 Related Works
2.1 Data Contamination
MLLMs are often pre-trained on massive, diverse datasets—often scraped from the web, which increases the risk of evaluation data overlapping (Dodge et al. 2021; Zhou et al. 2023). In the post-training phase, models are further fine-tuned on large human-annotated or synthetic datasets that may resemble evaluation tasks, further compounding contamination risks. Although retrieval-based detection methods (Golchin and Surdeanu 2023a, b; Yang et al. 2023) exist, the sheer scale and complexity of training corpora make it difficult to entirely exclude evaluation data. Additionally, many MLLMs keep their training data proprietary, complicating the accurate assessment of their true performance and highlighting the need for fair and reliable benchmarks. To address this issue, dynamic benchmarking has been proposed (Zhu et al. 2023; Fan et al. 2023; Lei et al. 2023). In this paper, we make a step forward in the safety dynamic benchmarking for MLLMs.
2.2 MLLM Safety Evaluation
Despite the great success of MLLMs in multimodal understanding and reasoning, their potential safety issues, such as truthfulness, value misalignment, and misuse, still pose significant challenges. Efforts have been made to evaluate the safety of MLLMs. Liu et al. (2023b) proposed MMsafetybench, a VQA dataset covering 13 harmful scenarios to assess MLLMs’ safety. Ch3ef (Shi et al. 2024) adopts ”Helpful, Honest, and Harmless” as safety evaluation criteria. Other benchmarks Hu et al. (2024); Gu et al. (2024) also investigate safety from different degrees. However, these benchmarks are manually constructed and lack updating. Their fixed complexity and diversity can’t match the fast progress of MLLMs. To address this issue, we propose SDEval to make dynamic evaluation, which injects randomness into existing data for MLLM safety benchmarking.
2.3 Dynamic Evaluation
Recently, the evaluations for MLLMs have gained much attention from both academia and industry (Yao 2025). However, existing evaluation benchmarks are static and have data leakage issues, so they can not match the fast progress of MLLMs. To solve this, some researchers have pioneered the exploration of dynamic evaluation. Zhu et al. propose DyVal V1 and V2 for the evaluation of LLMs (Zhu et al. 2023, 2024). Recently, Yang et al. (2024b) and Zhou et al. (2025) transfer this insight into the multi-modal domain and propose a vision-language bootstrapping strategy. However, these methods are not suitable for safety evaluation due to that their scenarios are limited to specific multiple-choice questions, while safety evaluation is usually open-ended and does not have fixed answers. Different from existing strategies, we construct a safety-centric dynamic framework, SDEval. Starting from any original benchmark, we can endlessly generate the variants with flexible complexity and lower contamination rates.
3 Method
In this section, we give a detailed introduction to our safety-centric dynamic evaluation framework, SDEval.
3.1 Overview
As shown in Figure 3, SDEval leverages multimodal dynamic strategies to modify the original benchmark, and then evaluates MLLM safety with the modified samples. Given a sample from the original benchmark, where represents the textual prompt and is the image, the dynamic generation process can be formulated as: , where represents the set of dynamic strategies and is the updated text-image pair.
SDEval generates new text-image pairs based on three dynamic mechanisms, including: a) Text dynamics, which aims to evaluate whether MLLMs have a robust understanding of the safety risks implied by different language expressions; b) Image dynamics, which aims to measure whether MLLMs can figure out the risk factors in the image; c) Text-image dynamics, which aims to test if MLLMs are influenced by cross-modality harmful contents. Additionally, in order to ensure the dynamically generated samples are semantically consistent with the original samples, we design a validator agent to verify them. Each modified sample should be validated to guarantee semantic consistency. These strategies, inspired by real-world safety concerns and jailbreaking tricks, pose a huge challenge to MLLMs. We use the new dynamically generated benchmark to evaluate MLLMs, and adopt a scorer to judge the harmfulness of model responses.
3.2 Text Dynamics
Language understanding is crucial for MLLMs, recent jailbreak research of MLLMs (Chao et al. 2023; Deng et al. 2023; Liu et al. 2023c) revealed that current MLLMs are sensitive to input texts. Similar to DME (Yang et al. 2024b), we construct text dynamics from the human-centric perspective–humans often adopt strategies such as replacing sensitive words, reorganizing sentences, or combining multiple languages (e.g., English and Chinese in one sentence) to form new sentences while maintaining semantics to circumvent safety review. Specifically, we utilize six dynamic strategies to modify the text prompt of sample :
Word Replacement
Given that some safety review mechanisms usually identify keywords, it is effective and reasonable to replace the words in the original sentence to perform text dynamic processing. Inspired by (Zhu et al. 2024), we prompt LLMs to replace no more than five words of each text prompt, , using synonyms or contextually similar words. For example, the word religious may be replaced by its synonyms like faith-based or faithful.
Sentence Paraphrasing
Inspired by the fact that humans may use different sentence structures to express the same meaning, we utilize sentence paraphrasing for text dynamics. This method centers on reframing questions while preserving their core concept. These rephrased questions test MLLMs’ ability to comprehend the question’s essence, moving beyond mere surface-level recognition.
Adding Descriptions
Following DME (Yang et al. 2024b), we utilize GPT-4o (OpenAI 2024) to add extra relevant/irrelevant descriptions into the original text, which may distract the model’s attention, thereby reducing its control over safety. Specifically, for adding relevant descriptions, we employ GPT-4o (OpenAI 2024) to analyze the image and generate a caption about the image, then add the caption in front of the original text; for irrelevant descriptions, we prompt GPT-4o (OpenAI 2024) to add descriptions that are not related to the image , and we append the irrelevant descriptions after the original text prompt.
Making Typos
Given the fact that humans often deliberately make spelling mistakes or repeat certain letters in specific words to evade safety review. These operations will not change the meaning of the sentence and will not affect the reader’s normal reading. Similar to (Vega et al. 2024), we utilize GPT-4o to make typos for each word in the given sentence by randomly selecting from the strategies of repeating, spelling mistakes, and special wrong characters.
Linguistic Mix
Considering the potential safety risks caused by the inaccurate recognition of multiple languages, we use a multilingual hybrid strategy to dynamically update text . Specifically, we prompt GPT-4o (OpenAI 2024) to reconstruct the original sentence from a single language into a combination of multiple languages, including Chinese, English, Russian, French, Japanese, and Korean.
Chain-of-Thought
We simply modified the question by adding the text: answer step by step to ask the MLLMs to answer the question in a chain-of-thought paradigm.
3.3 Image Dynamics
Harmfulness in image content is also critical for MLLM safety evaluation. Pioneering research (Yang et al. 2024b) proves that significant overlap exists between evaluation benchmark and MLLM training data, leading to data leakage problems. To address this, we adopt the strategy of dynamically updating and evaluating. The image strategies utilized can be divided into two parts:
Basic Augmentation
We utilize several basic augmentation operations to obtain new images:
1) Spatial Transformation: Inspired by (Zhou et al. 2025), we expand the original image boundaries by random padding, whose width is uniformly sampled from the range of the original image size. After that, random flipping is also adopted. This operation helps in testing MLLMs’ ability to recognize and locate harmful objects under different spatial transforms.
2) Color Transformation: The original image’s color scheme is flipped. Additionally, salt-and-pepper noise with randomly distributed density is injected. This process evaluates the model’s robustness against significant color shifts and visual disturbances.
Generation and Manipulation
Using a generative manner to obtain images that are different from the original ones can further reduce data leakage. We synthesize new samples via caption-guided generation and content manipulation. Furthermore, to make these synthesized images semantically consistent with the original ones, we adopt a validator to filter out inconsistent samples:
1) For generation, we prompt GPT-4o (OpenAI 2024) to generate a comprehensive caption, which highlights the layout, object details, and scene descriptions. Then, we guide Stable-Diffusion-3.5-Large (Esser et al. 2024) with the caption to generate a new version of the original image. Then, we leverage GPT-4o (OpenAI 2024) to verify whether the generated image maintains essential concepts of the original one. If not, we repeat this process until success.
2) For manipulation, we edit the original image via inserting objects, inserting texts, and style transfer. Specifically, we prompt GPT-4o (OpenAI 2024) to analyze the image and answer whether and how to conduct the following editing: (a) Inserting objects: Return the name of an object that can be inserted into the image without affecting the main content and safety of the image and the location that the object should be inserted; (b) Inserting texts: Return appropriate texts that can be inserted into the image without affecting semantics of the image. (c) Style transfer: Choose a suitable style from {Watercolor style, Sketch style, Comic style} based on the image. Then we conduct manipulation operations by utilizing the ICEdit (Zhang et al. 2025c) model. Examples are presented in Figure 4.
| ASD () | PAR () | |||||||||||||
| Model | Privacy | Bias | Toxicity | Truthfulness | Legality | Avg. | Vanilla | Privacy | Bias | Toxicity | Truthfulness | Legality | Avg. | Vanilla |
| GPT-4o | 40.74 | 28.57 | 31.80 | 26.96 | 35.83 | 32.78 (3.560) | 29.22 | 11.57 | 22.16 | 11.53 | 74.56 | 3.750 | 24.71 (15.67) | 40.38 |
| Gemini-2.5-Pro | 38.58 | 29.25 | 37.85 | 25.40 | 34.44 | 31.11 (2.240) | 28.87 | 32.87 | 45.19 | 13.83 | 83.65 | 37.08 | 42.52 (7.150) | 49.67 |
| Claude-4-Sonnet | 24.85 | 10.79 | 24.59 | 33.80 | 33.06 | 25.42 (1.930) | 23.49 | 53.24 | 74.64 | 30.84 | 62.40 | 38.33 | 51.89 (4.480) | 56.37 |
| o3 | 45.06 | 36.54 | 37.75 | 17.33 | 38.75 | 35.08 (2.420) | 32.66 | 26.94 | 33.41 | 11.53 | 61.02 | 18.30 | 30.24 (13.16) | 43.40 |
| LLaVA-V1.5-7B | 48.46 | 45.53 | 35.83 | 45.43 | 45.28 | 45.11(3.380) | 41.73 | 11.11 | 10.50 | 3.750 | 40.82 | 7.080 | 14.65 (4.150) | 18.80 |
| LLaVA-V1.5-13B | 49.54 | 49.56 | 35.64 | 37.19 | 43.75 | 43.14 (3.180) | 39.96 | 10.19 | 12.54 | 4.900 | 48.31 | 5.420 | 16.27 (3.960) | 20.23 |
| Qwen-VL-2.5-7B | 49.85 | 25.17 | 31.32 | 58.13 | 36.39 | 40.17 (13.71) | 29.46 | 17.13 | 44.02 | 20.75 | 70.79 | 17.08 | 33.96 (10.08) | 44.04 |
| Qwen-VL-2.5-72B | 39.35 | 25.07 | 30.84 | 28.55 | 39.72 | 32.76 (4.680) | 28.08 | 23.15 | 34.69 | 16.71 | 88.45 | 25.00 | 37.60 (8.730) | 46.33 |
| Qwen-VL-2-7B | 47.69 | 36.44 | 33.62 | 31.75 | 37.78 | 37.46 (5.430) | 32.03 | 16.67 | 27.99 | 15.56 | 61.42 | 17.08 | 27.74 (4.880) | 32.62 |
| Qwen-VL-2-72B | 50.31 | 30.52 | 35.25 | 17.40 | 38.61 | 34.42 (0.530) | 33.89 | 14.81 | 39.65 | 14.12 | 70.43 | 17.92 | 31.39 (4.280) | 35.67 |
| Yi-VL-6B | 41.51 | 47.62 | 34.68 | 38.65 | 39.86 | 40.46 (0.860) | 39.60 | 11.11 | 8.750 | 4.900 | 45.92 | 9.580 | 16.05 (3.110) | 19.16 |
| Yi-VL-34B | 41.67 | 43.83 | 35.46 | 32.56 | 38.33 | 38.41 (2.400) | 36.01 | 14.81 | 11.95 | 5.480 | 49.97 | 15.00 | 19.44 (2.780) | 22.22 |
| DeepSeek-VL | 43.52 | 39.36 | 35.64 | 67.34 | 35.56 | 44.31 (6.780) | 37.53 | 14.35 | 3.210 | 1.440 | 34.70 | 5.830 | 11.91 (10.78) | 22.69 |
| InternVL-Chat-V1.5 | 39.20 | 25.56 | 31.32 | 82.11 | 32.64 | 42.17 (9.760) | 32.41 | 18.06 | 38.19 | 11.53 | 60.78 | 17.92 | 29.30 (10.89) | 40.19 |
| InternVL-3-9B | 52.47 | 34.11 | 35.25 | 35.44 | 40.69 | 39.59 (6.310) | 33.28 | 13.43 | 31.78 | 12.39 | 61.06 | 12.50 | 26.23 (8.630) | 34.86 |
| InternVL-3-14B | 50.93 | 27.70 | 32.56 | 35.68 | 40.28 | 37.43 (5.350) | 32.08 | 14.35 | 41.98 | 18.16 | 67.28 | 12.92 | 30.94 (5.650) | 36.59 |
| InternVL-3-78B | 47.66 | 36.55 | 36.41 | 37.33 | 38.75 | 39.34 (9.240) | 30.04 | 3.240 | 13.41 | 12.10 | 73.97 | 2.500 | 21.40 (17.64) | 39.04 |
3.4 Text-Image Dynamics
To further improve the diversity of generation, we design text-image dynamic strategies to explore the cross-modality interaction. We integrate both text and image content to generate new text-image pairs, aiming to test if MLLMs are influenced by the combination of text-image dynamics. As shown in Figure 3, it mainly consists of two types, Text-to-Image and Image-to-Text generation. Additionally, we take into consideration the cross-modality jailbreaking tricks.
Text-to-Image Generation
Section 3.2 introduces a set of text dynamic strategies. In this text-to-image strategy, we aim to inject text perturbations into images via cross-modal generation, thereby impacting MLLM safety. This strategy amplifies the influence of text dynamics through cross-modal interactions. Specifically, we first sample a from text dynamics, then we feed the sampled and original image into GPT-4o to generate an image caption and extract the safety-related keywords, which will then be used as a prompt to generate new images via stable diffusion (Esser et al. 2024). We then utilize the sampled text and the generated image as the new text-image pair.
Image-to-Text Generation
In contrast to text-to-image strategies, we inject image perturbations into text prompts, thereby amplifying the impact of image dynamics on MLLM safety. This strategy aims to distract MLLMs by confounding perturbed content with original content. Specifically, we first sample a generated image from image dynamics, and then we feed the original text and sampled image into GPT-4o to obtain a safety-centric image caption. Finally, we prepend the generated caption in front of the original text as . and form the new pair.
Cross-modal Jailbreaking
We mainly investigate two jailbreaking tricks. a) Figstep (Gong et al. 2023) proves that MLLMs can recognize and answer the typographic questions in images. The safety guardrails of MLLMs are ineffective against the typographic visual prompts. Even if the LLM part of MLLMs has been safety aligned in advance, the visual inputs could introduce new risks since the visual embedding space is not safety aligned to the LLM’s embedding space. Based on this, we directly replace the text prompts with their typographic version. Thus, in this trick, only the typographic prompt and the original image are fed into MLLMs. b) HADES (Li et al. 2024a) empirically validates that MLLMs can be significantly affected by the unsafe words contained in images rather than texts, so we utilize GPT-4o to extract key information, which is strongly related to the safety content, and then inject it into the original image. This trick moves the unsafe content from the texts to the images without changing the original semantics.
| Models | #VR() | #DR() | #VW() | #DW() | #VS() | #DS() |
| GPT-4o | 53.01 | 48.33 | 10.22 | 4.510 | 58.50 | 52.83 (5.670) |
| Gemini2.5-Pro | 3.260 | 5.890 | 34.90 | 24.05 | 38.15 | 29.94 (8.210) |
| Claude-4-Sonnet | 11.83 | 18.83 | 35.52 | 25.44 | 47.34 | 44.27 (3.070) |
| o3 | 46.18 | 43.66 | 13.21 | 11.43 | 59.39 | 55.09 (4.300) |
| LLaVA-V1.5-7B | 0.000 | 0.090 | 6.600 | 3.210 | 6.600 | 3.300 (3.000) |
| LLaVA-V1.5-13B | 0.000 | 0.220 | 8.650 | 4.060 | 8.650 | 4.280 (4.170) |
| Qwen-VL-2-7B | 1.470 | 2.810 | 8.520 | 4.600 | 10.00 | 7.410 (2.590) |
| Qwen-VL-2-72B | 2.230 | 1.980 | 11.56 | 8.440 | 13.79 | 10.42 (3.370) |
| Qwen-VL-2.5-7B | 0.760 | 0.040 | 9.100 | 9.370 | 9.860 | 9.410 (0.450) |
| Qwen-VL-2.5-72B | 0.270 | 2.230 | 14.41 | 11.56 | 14.68 | 13.79 (0.890) |
| Yi-VL-6B | 0.130 | 0.110 | 6.110 | 4.270 | 6.250 | 4.400 (1.850) |
| Yi-VL-34B | 0.220 | 0.580 | 6.740 | 3.440 | 6.960 | 4.020 (2.940) |
| DeepSeek-VL | 2.370 | 3.440 | 4.960 | 0.760 | 7.360 | 4.190 (3.170) |
| InternVL-Chat-V1.5 | 0.400 | 3.170 | 10.93 | 7.810 | 11.33 | 10.98 (0.350) |
| InternVL-3-9B | 0.620 | 0.670 | 8.300 | 5.350 | 8.920 | 6.020 (2.900) |
| InternVL-3-14B | 1.740 | 1.430 | 10.17 | 6.290 | 11.91 | 7.720 (4.190) |
| InternVL-3-78B | 2.230 | 1.080 | 11.56 | 7.440 | 13.79 | 8.520(5.270) |
4 Experiment
4.1 Experimental Setup
Dataset
In order to evaluate the effectiveness of SDEval, we select two comprehensive multimodal large language models safety benchmark MLLMGuard (Gu et al. 2024) and VLSBench (Hu et al. 2024) for dynamic evaluation.
MLLMguard (Gu et al. 2024) develops a collection of adversarial examples to test the ability of MLLMs to identify and counteract attacks orchestrated by red teams. We follow (Gu et al. 2024) to evaluate the proposed two metrics, Attack Success Degree (ASD), which quantifies the extent to which the responses are harmless, and Perfect Answer Rate (PAR), which represents the proportion of safe and responsible responses among all responses.
VLSBench (Hu et al. 2024) aims to address the existing issues in current multimodal safety benchmarks, called Visual Safety Information Leakage (VSIL). Following VLSBench, we calculate the Safety Rate (SR) by considering the total number of safe refusals and safe warnings.
Evaluated MLLMs
We evaluate four close source MLLMs: GPT-4o (OpenAI 2024), o3 (OpenAI 2025), Claude-4-Sonnet (Anthropic. 2025), and Gemini2.5-Pro (Gemini Team 2025), and extensive open-sourced models: Qwen-VL family (Bai et al. 2025): Qwen2.5-VL-7B, Qwen2.5-VL-72B, Qwen2-VL-7B, Qwen2-VL-72B; Yi-VL family (Young et al. 2024): Yi-VI-6B, Yi-VI-34B; InternVL family (Zhu et al. 2025): InternVL-3-9B, InternVL-3-14B, InternVL-Chat-V1-5; LLaVA family (Liu et al. 2023a): LLaVA-V1.5-7B, LLaVA-V1.5-13B. To ensure a standardized comparison, we set the generation temperature to 0 for all models. All experiments are conducted on 4 NVIDIA A800 GPUs.
4.2 Results of Dynamic Evaluation
To understand the impact of each strategy, we conduct comprehensive experiments to figure out the most powerful dynamic operations in Section 4.3. We select the most influential strategies (Word Replacement and Figstep) to conduct experiments on MLLMGuard (Gu et al. 2024) and VLSBench (Hu et al. 2024).
Results on MLLMGuard
As shown in Table 1, we exploit the mentioned dynamic strategies to conduct experiments. The results show that the dynamic strategies significantly reduce the safety rate of MLLMs, indicating that the degree of safety control of each model is easily disturbed. Compared with ASD, the PAR index is reduced more after applying dynamics, which means that the dynamic strategies improve the difficulty and complexity of the original data. Thus, the model’s attention is distracted by the dynamic strategy, and the control over safety is weakened.
| Strategy | Variants | ASD() | PAR() | ||
| Oringinal | Vanilla | 32.21 | 40.19 | ||
Word Replacement | 38.71 (6.300) | 26.94 (13.25) | ||
| Sentence Paraphrasing | 36.68 (4.270) | 26.07 (14.12) | |||
| Adding Descriptions | 32.97 (0.560) | 30.67 (9.520) | |||
| Making Typos | 35.56 (3.150) | 30.70 (9.490) | |||
| Linguistic Mix | 33.23 (0.820) | 32.95 (7.240) | |||
| Chain-of-Thought | 32.67 (0.260) | 30.96 (9.230) | |||
Adding Texts | 33.73 (1.320) | 29.58 (10.61) | ||
| Adding Objects | 39.41 (7.000) | 26.45(13.74) | |||
| Generation | 34.89 (2.480) | 26.14 (14.05) | |||
| Augmentation | 33.97 (1.560) | 36.46 (3.730) | |||
| Style Transfer | 35.20 (2.790) | 26.38 (13.81) | |||
Text-to-Image | 35.10 (2.690) | 24.36 (15.83) | ||
| Image-to-Text | 34.89 (2.480) | 31.65 (8.540) | |||
| FigStep | 41.96 (9.550) | 17.08 (23.11) | |||
| HADES | 35.71 (3.500) | 28.54 (11.65) |
Results on VLSBench
As shown in Table 2, the closed-source model performs much better than open-source models in both original and dynamic benchmarking. Claude-4-Sonnet (Anthropic. 2025) gets the highest safety rate, and it is also safer than others. Additionally, after the dynamic strategies, the safety of all MLLMs decreases, and the proportion of answers judged as containing warnings decreases more, which shows that the dynamic strategy causes more safety risks for MLLMs, thereby reducing the safety rate of the model. The dynamic evaluation results show that the MLLMs still face huge safety risks.
Can MLLMs Cope With Safety Dynamic Evaluation Well?
Safety dynamic evaluation aims at mitigating data leakage and making the evaluation framework dynamically scalable. After applying the proposed dynamic strategies, the safety performance of all MLLMs has been greatly reduced, which indicates that MLLMs may just memory the safe answers and unsafe answers, and they do not really understand the unsafe factors. The performance decrease also shows that the data leakage issue in the current safety benchmarks is significantly alleviated.
If the Scaling Law Still Works for Safety Dynamic Evaluation?
We can conclude from the experiments that MLLMs with different parameters have different robustness for the same dynamic strategy and have no obvious correlation with the scaling law. An increase in model parameters does not significantly enhance safety levels across all dimensions, even leading to a drop in some cases (e.g., InternVL Family). We believe that the scale of parameters increases the performance of models, making them more effective at understanding human requirements and thus more likely to execute human input instructions, even if they are harmful.
Overall, these results show that current MLLMs are not good enough to cope with safety dynamic evaluation, suggesting there is data leakage in the current model training process, and current MLLMs still can’t handle safety issues well. How to ensure that the model’s safety and performance can develop in a balanced manner under the AI Law (Yang et al. 2024a) roadmap is still a huge challenge.
4.3 Ablation Study
| Model | MMVet | MMBench | ||
| Vanilla | Dynamic | Vanilla | Dynamic | |
| GPT-4o | 68.8 | 67.5 (1.30) | 83.4 | 81.8 (1.60) |
| o3 | 71.5 | 69.2 (2.30) | 84.8 | 82.9 (1.90) |
| Claude4-sonnet | 65.4 | 63.5 (1.90) | 86.8 | 83.6 (3.20) |
| Gemini2.5-Pro | 78.1 | 76.5 (1.60) | 90.1 | 87.5 (2.60) |
| LLaVA-v1.5-7B | 40.4 | 37.8 (2.60) | 66.5 | 64.0 (2.50) |
| LLaVA-v1.5-13B | 40.2 | 38.5 (1.70) | 69.2 | 65.8 (3.40) |
| Qwen2VL-72B | 74.0 | 70.8 (3.20) | 86.5 | 83.6 (2.90) |
| Qwen2VL-7B | 62.0 | 57.3 (4.70) | 83.0 | 80.4(2.60) |
| Qwen2.5VL-72B | 76.2 | 72.3 (3.90) | 88.6 | 86.4 (2.20) |
| Qwen2.5VL-7B | 67.1 | 63.9 (3.20) | 83.5 | 79.3 (4.20) |
| InternVL3-9B | 76.2 | 72.9 (3.30) | 83.4 | 81.8 (1.60) |
| InternVL3-14B | 80.2 | 77.9 (2.30) | 85.6 | 82.8 (2.80) |
| InternVL-Chat-V1.5 | 61.5 | 58.1 (3.40) | 82.2 | 80.5 (1.70) |
| InternVL3-78B | 81.3 | 78.6 (2.70) | 89.0 | 87.6 (2.40) |
| Yi-VL-6B | 28.0 | 26.8 (1.20) | 68.4 | 65.8 (2.60) |
| Yi-VL-34B | 30.5 | 26.3 (4.20) | 72.4 | 70.7 (1.70) |
| DeepSeek-VL | 41.5 | 34.5 (7.00) | 84.1 | 82.1 (2.00) |
In order to explore the impact of each dynamic strategy on safety evaluation, we select some powerful open-source multimodal large language models, including Qwen-VL family (Bai et al. 2025), InternVL family (Zhu et al. 2025), Yi-VL family (Young et al. 2024), to conduct comprehensive ablation experiments for each dynamic operation based on MLLMGuard (Gu et al. 2024) Benchmark. We only show the results of InternVL-Chat-V1.5 in Table 3. More ablation experiment results can be found in our Appendix.
Text Dynamics
We adopt the dynamic strategies mentioned in Section 3.2 to evaluate the MLLMs. As can be seen in Table 3, we find that all dynamic strategies reduced the safety performance of the model. Among them, the Chain-of-Thought dynamic strategy has the least impact on the safety performance of the model. We attributed this to the fact that the Chain-of-Thought strategy only encourages the model to analyze the problem step by step, which does not significantly increase the safety risk of the model. However, the word replacement strategy caused a significant decrease in perfect answer rate and the greatest increase in attack success degree. We believe that this is because the semantic expression deceived the model due to the replacement of some words, causing the model to output unsafe content.
Image Dynamics
Also, whether the image content is safe is very important for MLLM’s safety. As we presented in the Table 3, all image dynamic strategies have caused varying degrees of degradation in the safety performance of the model. We find that the strategy of add object risks the safety of MLLMs most, while the add text strategy has the least impact on the model’s safety. This means that add object will distract the model and thus prevent MLLMs from capturing critical safety content relevant to the problem. While the strategy of add text has less impact on the MLLM’s understanding of images.
Text-Image Dynamics
As shown in the Table 3, when applying text-image dynamic strategies, the safety performance also falls sharply, which means that combining images and text can easily make MLLM jailbreak, thereby reducing safety performance. Among the proposed strategies, Figstep results in the biggest increase in the attack success degree, which means that combining images and text in a way similar to Figstep will pose a huge threat to the safety performance of the model. Current models are still struggling to cope well with jailbreak attacks.
5 Safety-Capability Balance
Current MLLMs achieve remarkable progress in intelligent capability, while they may fall short in safety, which may cause large risks for the entire society. How to ensure “intelligence for good” and achieving balanced development of capabilities and safety is an important issue that must be paid attention to and resolved. In this section, we further conduct experiments on MLLM capability evaluation to figure out the dynamic evaluation for safety and capability.
Results on MLLM Capability Evaluation
As SDEval is a general framework, it can also be utilized for MLLM capability evaluation. The input for the capability evaluation of MLLM still consists of two parts: text and images. We can utilize the proposed text, image, and text-image dynamic strategies to conduct capability evaluation for MLLM, which can relieve the data leakage issues. And here we choose the same dynamic strategies as safety dynamic evaluation for capability evaluation. We apply SDEval dynamic evaluation on some popular benchmark datasets (MMVet (Yu et al. 2023), MMBench (Liu et al. 2023d)) to assess current MLLMs. As can be seen in Table 4, after applying the proposed dynamic strategies, all the metrics on MMVet and MMBench decreased, which means that our SDEval framework is not only suited for safety evaluation, but also suited for more general evaluation benchmarks. More results can be seen in our Appendix.
(a) Ranking
(b) Dynamic Change
Balance between Safety and Capability
AI theory (Zhang et al. 2025b; Yang et al. 2024a) hypothesizes that the development of AGI should consider the balance of model performance and safety. The safety and capability of AI are generally balanced along a roadway. In the short term, rotation is allowed, but in the long term, it should not be lower than , as in the current state, or higher than , which would hinder development and industrial application. We weighted the safety and capability scores of the selected datasets after dynamic evaluation according to the dataset size and drew a capability-safety scatter plot based on this. As can be seen in Figure 5 (a), we present the weighted ranking figure: Claude-4-Sonnet outperforms all the models on safety, and it also has a good performance in the intelligent capability, and Gemini-2.5-Pro has achieved an excellent balance between safety and capability. As presented in Figure 5 (b), most MLLMs have worse robustness in safety, resulting in more significant safety performance loss when performing dynamic strategies, which highlights the need to strengthen the model’s safety ability in future development.
6 Conclusion
In this paper, we propose SDEval, a MLLM safety dynamic evaluation framework to mitigate data leakage and the static complexity issues. SDEval introduces a comprehensive cross-modal dynamic evaluation framework, incorporating diverse text, image, and text-image dynamic strategies, which generate new samples from original benchmarks to test model safety. Experimental results demonstrate that our approach effectively mitigates the data leakage problem and enhances the complexity of static datasets, enabling benchmarks to co-evolve with models. Furthermore, SDEval’s versatility allows its application to various existing MLLM safety benchmarks. Through extensive evaluations, we uncover safety risks in current MLLMs, highlighting areas for potential improvement of MLLM safety.